OpenStudy (dls):

For a reversible reaction taking place at 127 degree celsius and 1ATM pressure.delta G=23.5kj/mol. for the forward reaction,calculate value of eqm.constant for the reaction

OpenStudy (dls):

\[\LARGE \delta G=-2.303RTlogK\] I used this,getting wrong answer.

OpenStudy (dls):

\[\LARGE 23500=-2.303 \times 8.314 \times 400 \times logK\]

OpenStudy (yrelhan4):

relation to yahi hota hai mere khayaal.

OpenStudy (dls):

\[\LARGE K=10^{0.30}\] gettting something likethis

OpenStudy (dls):

not sure if i substituted correctly,the values

OpenStudy (yrelhan4):

127 degree Celsius hai. convert it into kelvin

OpenStudy (yrelhan4):

acha kiya hai. thik hai.

OpenStudy (dls):

did,400K

OpenStudy (yrelhan4):

e^7.07 hai answer?

OpenStudy (dls):

arre numerically nikaalo

OpenStudy (dls):

e wagera nahi :/

OpenStudy (yrelhan4):

1176.14

OpenStudy (dls):

its 8.59*10^-4

OpenStudy (yrelhan4):

dekhta hun. get back to you later.

OpenStudy (dls):

dont do it after 12AM :P

OpenStudy (yrelhan4):

aa to rha hai. kya problem hai? calculator se krna pdega.

OpenStudy (dls):

kya aya logk?

OpenStudy (dls):

and meri substituted values sahi thi? yuhuuuuuuuuuuu @yrelhan4

OpenStudy (yrelhan4):

haan haan thik thi. mereko yaad nhi kya aaya tha. pehle log(1/K) niklega. kyun ki -logK=log(1/K), fir 10^(---)=abc bhi krke 1/K niklega, fir 1/abc krke K niklega.

OpenStudy (dls):

okay :|

OpenStudy (dls):

TY

OpenStudy (yrelhan4):

yw. :/

OpenStudy (shubhamsrg):

it'll be -23500 = -RT lnK

OpenStudy (yrelhan4):

why -23500?

OpenStudy (shubhamsrg):

forward reacn would imply dG < 0

OpenStudy (yrelhan4):

its a reversible reaction.

OpenStudy (shubhamsrg):

question wants us to calc K for forward reacn

OpenStudy (yrelhan4):

and delta G for the forward reaction is given. if it was negative it would have been given

OpenStudy (shubhamsrg):

dG for forward reacn ? o.O It means nothing. dG for the whole reacn is given. for it to proceed forward, dG< 0 .

OpenStudy (yrelhan4):

hmm. :O but i still have my doubts. when you say deltaG=-nFE(electrochem), and you are given a deltaG for a reaction, and you are a deltaG, do you do -deltaG then too? i think the negative sign on the RHS takes care of that. i dont know. i may be wrong.

OpenStudy (shubhamsrg):

Well anytime dG is put with the sign. The sign is always given in the ques. This was a different ques which asks us to calc K for forward reacn, Well I don't understand the ques practically, but theoretically, you will have to take dG<0. I have cross checked the ans via that.

OpenStudy (yrelhan4):

-23500 = -RT lnK gives you 8.59*10^-4?

OpenStudy (shubhamsrg):

dafaq! it doesn't! o.O

OpenStudy (shubhamsrg):

I was wrong1 -_-

OpenStudy (shubhamsrg):

got carried away, so sorry! :|

OpenStudy (yrelhan4):

haha.. np. :P